OpenStudy (elijahrich1113):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3? a. -21 b. -16 c. -10 d. -1

OpenStudy (plainntall):

What is an inverse function? Isn't that when we switch the x and y? y=4(x+3)-5=4x+12-5=4x+7 switch x with y x=4y+7 then solve for y?

OpenStudy (elijahrich1113):

Yes kind of